1. Outback Outfitters sells recreational equipment. One of the company’s products, a small camp stove, sells for $100 per unit. Variable expenses are $70 per stove, and fixed expenses associated with the stove total $150,000 per month.

Required:

1. What is the break-even point in unit sales and in dollar sales?

2. If the variable expenses per stove increase as a percentage of the selling price, will it result in a higher or a lower break-even point? (Assume that the fixed expenses remain unchanged.)

3. At present, the company is selling 11,000 stoves per month. The sales manager is convinced that a 10% reduction in the selling price would result in a 25% increase in monthly sales of stoves. Prepare two contribution format income statements, one under present operating conditions, and one as operations would appear after the proposed changes.

4. Refer to the data in Required 3. How many stoves would have to be sold at the new selling price to attain a target profit of $71,000 per month?

2. Northwood Company manufactures basketballs. The company has a ball that sells for $25. At present, the ball is manufactured in a small plant that relies heavily on direct labor workers. Thus, variable expenses are high, totaling $15.00 per ball, of which 60% is direct labor cost.

Last year, the company sold 30,000 of these balls, with the following results:

Sales (30,000 balls) $ 750,000
Variable expenses 450,000
Contribution margin 300,000
Fixed expenses 210,000
Net operating income $ 90,000

Required:

1. Compute (a) last year's CM ratio and the break-even point in balls, and (b) the degree of operating leverage at last year’s sales level.

2. Due to an increase in labor rates, the company estimates that next year's variable expenses will increase by $3.00 per ball. If this change takes place and the selling price per ball remains constant at $25.00, what will be next year's CM ratio and the break-even point in balls?

3. Refer to the data in (2) above. If the expected change in variable expenses takes place, how many balls will have to be sold next year to earn the same net operating income, $90,000, as last year?

4. Refer again to the data in (2) above. The president feels that the company must raise the selling price of its basketballs. If Northwood Company wants to maintain the same CM ratio as last year (as computed in requirement 1a), what selling price per ball must it charge next year to cover the increased labor costs?

5. Refer to the original data. The company is discussing the construction of a new, automated manufacturing plant. The new plant would slash variable expenses per ball by 40.00%, but it would cause fixed expenses per year to double. If the new plant is built, what would be the company’s new CM ratio and new break-even point in balls?

6. Refer to the data in (5) above.

a. If the new plant is built, how many balls will have to be sold next year to earn the same net operating income, $90,000, as last year?

b. Assume the new plant is built and that next year the company manufactures and sells 30,000 balls (the same number as sold last year). Prepare a contribution format income statement and Compute the degree of operating leverage.
Answer #1
Please give positive ratings so I can keep answering. It would help me a lot. Please comment if you have any query. Thanks!
As per HOMEWORKLIB POLICY I can only solve first four questions. So you have tio raise new question for Northwood. So sorry for any inconvience. Thanks!
Outback Outfitters
Answer 1 Amount $ Note
Sell price              100.00 E
Less: Variable expenses                70.00
Contribution per unit                30.00 A
Fixed expense      150,000.00 B
Break even units          5,000.00 C=B/A
Break even $      500,000.00 D=C*E
Answer 2
If variable cost increase in same % as sales then CVP ratio or contribution per unit will also increase in same % as sales so break even will decrease.
So higher breakeven point.
//img.homeworklib.com/questions/35c75990-79a4-11ea-854d-cf4e6376cf66.gif?x-oss-process=image/resize,w_560
Answer 3
Current Sell price              100.00
Reduction by 10%                10.00
Revised Sell price                90.00
//img.homeworklib.com/questions/35c75990-79a4-11ea-854d-cf4e6376cf66.gif?x-oss-process=image/resize,w_560
Current Sales units        11,000.00
Increase by 25%          2,750.00
Revised Sales units        13,750.00
Income statement Present Proposed
11000 Stoves 13750 Stoves Note
Total Per Unit Total Per Unit
//img.homeworklib.com/questions/35c75990-79a4-11ea-854d-cf4e6376cf66.gif?x-oss-process=image/resize,w_560
Sales
1,100,000.00     100.00 1,237,500.00        90.00
Les: Variable expenses      770,000.00        70.00      962,500.00        70.00
Contribution per unit      330,000.00        30.00      275,000.00        20.00 F
Fixed expense      150,000.00      150,000.00
Operating Income      180,000.00      125,000.00
//img.homeworklib.com/questions/35c75990-79a4-11ea-854d-cf4e6376cf66.gif?x-oss-process=image/resize,w_560
Decrease by
       55,000.00
Answer 4 Note
Target Profit        71,000.00
Add: Fixed expense      150,000.00
Target Contribution      221,000.00 G
Contribution per unit                20.00 F
Units to be sold        11,050.00 H=G/F